write the equation of the line passing through each of the given pairs of points. Write your results in slope intercept form where possible.
#20 is: (-1,3) and (4,-2)
#24 is: (2,-3) and (2,4)
Thanks!

Answer by checkley75(3666)

20) SLOPE=(-2-3)/(4+1)=-5/5=-1
NOW SUBSTITUTE 1 SET OF POINTS & SOLVE FOR THE Y INTERCEPT.
3=-1*1+b
3=-1+b
b=3+1
b=4
Y=X+4 ANSWER.
----------------------------------------------------------
24) SLOPE=(4+3)/(2-2)=7/0 THIS REPRESENTS A VERTICLE LINE THROUGH X=2
THUS THE EQUATION IS X=2.